OpenStudy (anonymous):

(x^2/25)-[(y-2)^2/4]=1 Find the foci and vertices

OpenStudy (anonymous):

Since it is a Hyperbola c^2 = a^2 + b^2 where c = Focii

Directrix (directrix):

@Yahoo! We're practicing (study session) for a supervised "in the flesh" test. So, it is okay to provide more details if you like. Just saying.

OpenStudy (anonymous):

I have my answers, but I'm not sure if it's correct

OpenStudy (anonymous):

Ok...Then 1st show ur Work @sakigirl

Directrix (directrix):

I guess the others left. V(5,2) and V(-5,2) F(-√29,2) and (√29, 2) Did anybody find the asymtotes?

OpenStudy (anonymous):

I get the same foci, but different vertices

Directrix (directrix):

Did you agree with the x part of the vertices?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What do you mean by y= 0 or 4?

Directrix (directrix):

Let me see your work.

OpenStudy (anonymous):

My work on how I found my vertices?

OpenStudy (anonymous):

I don't really do any work, I just look at it, and write them down haha

Directrix (directrix):

Well, what did you get?

Directrix (directrix):

Here is some good info on hyperbolas.

Directrix (directrix):

http://www.purplemath.com/modules/hyperbola2.htm

OpenStudy (anonymous):

V(5,4) and V(-5,4)

OpenStudy (anonymous):

How do you get 2? Even if you sq rt 4, you end up getting 2 or -6

Directrix (directrix):

Do you agree that on this hyperbola that the y-coordinates of the vertices and foci should be the same number, whatever that number may be?

OpenStudy (anonymous):

Oh wait, I understand

OpenStudy (anonymous):

I got mixed up. I needed to find the center to get it right

Directrix (directrix):

The center is at (0,2). Agree?

OpenStudy (anonymous):

Yes